OpenStudy (anonymous):

Please help!! I need help on 2 word problems!

OpenStudy (anonymous):

6.02] Johnston High School has a total of 82 boys and girls who play sports. If the number of boys is 16 more than twice the number of girls, how many boys play sports at this high school? 22 42 51 60

OpenStudy (anonymous):

What grade are you in?

OpenStudy (anonymous):

9th grade, u?

OpenStudy (anonymous):

4th. the answer in 60.

OpenStudy (nubeer):

no. of girls =x no. of boys are 16 more then twice the number og girls , so i think boys = 2x+16 total = no. of girls+ no. of boys.. you can find the value of x

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

2. x: girls, y: boys x+y=82 2x+16=y x+(2x+16)=82 3x+16=82 x=22 2x+16=y 2(22)+16=y 44+16=y 60=y There are 22 girls, and 60 boys.

OpenStudy (nubeer):

it's all the information given in the question.. it's all about english and understanding it.

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

Yep! Totally, got it! :D

OpenStudy (anonymous):

Best response?

OpenStudy (anonymous):

 Thx!